# Natural Language Generation from Visual Events: Challenges and Future Directions

### v2

@@ 2.4 Off-the-shelf Pre-trained VLMs
Qwen2.5-VL Bai et al. (2025) ViT Linear Qwen2.5
DeepSeek-VL Lu et al. (2024) SigLIP, SAM-B Linear (#2) DeepSeek
LLaVA-Next Liu et al. (2023) CLIP-ViT-L Linear Mistral
Video-LLaMA Zhang, Li, and Bing (2023) ViT-G/14 Q-Former LLAMA
m-PLUG-Owl3 Ye et al. (2025) SigLIP SO (400M) Linear Qwen2
Video-LLaVA Lin et al. (2024) OpenCLIP ViT-L Linear (#2) Vicuna v1.5
@@
Table 2: A selection of pre-trained multi-purpose VLMs that are used off-the-shelf for various NLG tasks involving visual events.
3 Evaluation
@@
Automatically generated text is notoriously difficult to evaluate. Most of the methods currently used to evaluate text describing visual events are borrowed from simpler NLG scenarios, and hence are not sufficient to fully understand the quality of language generated from visual input unfolding over time. These methods range from using traditional n-gram matching metrics to obtaining human judgments and ratings to, more recently, using off-the-shelf pre-trained VLMs assessing the generated output. In this section, we broadly classify evaluation methods into two main categories—automatic and human evaluation—and discuss several metrics widely used for vision-to-text tasks, along with the rationales for using them. Through this discussion, we also underline how the current approaches are lacking when it comes to assessing models in terms of the particular challenges relevant to the problem of language generation from visual events.
3.1 Automatic Evaluation
@@
Very often, tasks involving language generation from visual events assess a model-generated candidate text by comparing it to corresponding human-written references. Specifically, traditional metrics that were originally designed for evaluating machine translation and text summarization tasks—BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004)—are used to measure precision and recall of overlapping $\mathit{n}$-grams between the candidate and a reference text. Usually, metrics such as CIDEr Vedantam, Darrell, and Parikh (2015) and SPICE Anderson et al. (2016), which have been specifically developed for the evaluation of image and video captioning, are also used in conjunction with the above three metrics.
All the above-mentioned metrics rely on direct raw text comparisons of ground-truth references and model outputs. As this ground-truth data might not be always available, embedding-level reference-based evaluation metrics such as WMD Kusner et al. (2015), BERTScore Zhang* et al. (2020), and ViLBERTScore Lee et al. (2020) have been proposed. Recent work for Video Captioning, Video QA, and Movie Auto AD have used these metrics to measure the similarity between candidate and reference embeddings, obtained by projecting corresponding text into a common pre-trained semantic space Lin et al. (2022); Han et al. (2023a, 2024).
@@
Comparing model-generated text to ground-truth references, typically provided by crowd-workers or scraped from the internet, has various limitations when applied to the context of visually-grounded language generation. First, most reference-based metrics do not account for the visual modality upon which the generated text is conditioned. Second, reference-based metrics are generally inappropriate to accurately evaluate language that is not merely descriptive, but may involve creative or narrative concepts. This is the case for various tasks involving visual events, such as Video Captioning, Movie Auto AD, and Visual Storytelling.
On the language-only level, various reference-free metrics—not relying on any reference text—such as MAUVE Pillutla et al. (2021) and UNION Guan and Huang (2020) have been proposed. However, for visually-conditioned text generation tasks, this shift is still relatively recent, with a persistent emphasis on reference-based $\mathit{n}$-gram metrics to date. That said, various reference-free metrics have been recently proposed to assess different aspects of evaluation that are important for generation tasks that specifically concern visual events. For instance, metrics such as CLIPScore Hessel et al. (2021) and GROOViST Surikuchi, Pezzelle, and Fernández (2023) have been developed for evaluating visual grounding—the degree of alignment between the generated text and the visual input—in Video Captioning and Visual Storytelling tasks. Similarly, the RoViST Wang, Han, and Poon (2022) suite of metrics has been proposed to assess coherence, the extent of repetition, and visual grounding of the generated text. Yet other metrics, such as CRITIC Han et al. (2024) and ‘character matching’ Hong et al. (2023), are designed to evaluate task-specific aspects such as the accuracy of referencing to characters in the model outputs.
Finally, there is currently an increasing adoption and reliance on using pre-trained LLMs and VLMs as judges Li et al. (2024a). Essentially, these pre-trained general-purpose models are prompted to score or rate a model-generated response along any of the evaluation dimensions of interest, such as fluency or relevance (e.g., ‘How fluent is the generated text on a scale of 1 to 5?’). However, the effectiveness and reliability of this approach are still debated Bavaresco et al. (2024).
3.2 Human Evaluation
@@
Participants of these evaluation efforts are provided with a set of task-specific rubrics along with representative examples required for judging the model outputs. They are asked to assess the overall quality of model-generated outputs either independently (per sample) Surikuchi, Pezzelle, and Fernández (2023) or relative to outputs from other models Wang, Han, and Poon (2022). Alternatively, evaluators might be required to provide ratings for various criteria ranging from broad (e.g., text conciseness, fluency, grammatical correctness) to specific (e.g., factuality, hallucinations, expressiveness). The obtained scores are usually compared pairwise to rank models appropriately.
Some pre-trained VLM frameworks, such as LLaVA-RLHF Sun et al. (2024) leverage this qualitative feedback to optimize model parameters for learning to generate human-preferred text. Although human evaluation is still indispensable for several tasks, it is also expensive, time-consuming, and challenging. Defining clear evaluation protocols for ensuring the reliability and quality of human judgments is an active research area Kasai et al. (2022); Ruan et al. (2024).
@@
4 Key Challenges in Language Generation from Visual Events
As mentioned in the Introduction, tasks involving language generation from visual events present a unique set of challenges. These involve modeling the intricate relationship between temporally ordered visual events and the structure, content, and function of the language used to interpret, describe, or narrate them. We argue that solving these tasks requires models to be capable of identifying and managing such intricacies, leveraging the temporal, causal, and discourse-level patterns inherent in both visual data and language, as well as the many ways in which these aspects interact. In this section, we identify and describe some of these challenges and discuss how they are key to various tasks.
@@ 4 Key Challenges in Language Generation from Visual Events
Identifying and tracking entities is an important requisite for accurately interpreting actions and relationships between them. This has been underlined by several works in both language understanding Paperno et al. (2016); Kim and Schuster (2023) and computer vision domains Narayan et al. (2017); Luo et al. (2021). In our multimodal tasks, the availability of input signals from two modalities makes the aspect of disambiguating entities more challenging. For instance, in the Movie Auto AD or Visual Storytelling tasks where the visual input is heterogeneous, entities in the visual input tend to ‘disappear’ in some of the images or video frames at the intermediate temporal positions, while still being actively referenced in the textual input at the corresponding positions. Surikuchi, Pezzelle, and Fernández (2023) have denoted such cases as being temporally misaligned. To track entities accurately under such temporal misalignment, it is necessary to not only learn causal relations between people and objects in each of the modalities at corresponding positions, but also to obtain a cross-modal, cross-temporal representation of all the relationships relevant to the scene and the overarching narrative. For the Change Captioning, Video Captioning, or Video QA tasks, in which the input images or video frames are typically similar to each other, it is crucial to differentiate meaningful semantic entities and their changes from various distractions. While in Change Captioning, viewpoint changes or illuminations are considered as distractors and discarded, in the Video QA task, entities relevant to answering the question need to be differentiated from others for accurate tracking. Effective tracking of entities would therefore require accounting for changes in appearance (including disappearance), capturing interactions, and correctly identifying occlusions.
@@
Humans acquire language through perception and interaction with the environment Barsalou (2008); Iverson and Goldin-Meadow (2005) and consequently, this enables them to seamlessly ground language in visual data. Over the years, a great deal of work has been proposed to adapt the architecture and learning process of vision-language models for acquiring visual grounding Suglia, Konstas, and Lemon (2024). However, there are still significant challenges for achieving human-levels of grounding using computational models, and this becomes more apparent when looked at from the perspective of the tasks under investigation here. For instance, in Visual Storytelling or Movie Auto AD, the language used is often creative or abstract, including various sources of underspecification, such as context-dependent adverbs like ‘there’, ‘often’, or ‘today’ (’a photographer over there clicks…’). Recently, vision-language models have been shown to struggle with underspecified language, due to the challenges related to grounding underspecified terms to regions in the visual content Pezzelle (2023). Moreover, the amount of language informativeness—i.e., the degree of information required for identifying the correct object Coppock et al. (2020)—could be inadequate in tasks such as Video QA, particularly with the presence of confounding entities in various frames (e.g., input video of a football match and a question: ‘What is the color of the card the referee is holding?’). In addition, when grounding objects, models have often struggled to reliably capture spatial relationships Kamath, Hessel, and Chang (2023). To summarize, the language used in these tasks might not be merely descriptive but rather complementary to the content depicted in the corresponding images or videos. This makes visual grounding challenging without access to relevant additional external knowledge.
@@
To meet the communicative specificity of each task dealing with visual events, models would need to use additional information beyond what is available in the visual input. When performing Video Captioning or Video QA on specific domains such as news, the input video might not contain all the aspects needed to correctly describe its contents or answer questions about it Whitehead et al. (2018); Jin et al. (2023). To address this, various approaches often rely on using large pre-trained general purpose models or external knowledge bases such as ConceptNet Speer, Chin, and Havasi (2017) to retrieve both factual and commonsense information. This method is commonly referred to as retrieval-augmented generation (RAG) Lewis et al. (2020). Besides being sources for missing information, external knowledge bases are also leveraged for enriching the generated text with social or cultural contexts. For instance, in Visual Storytelling, some approaches use recognized visual objects in the input to retrieve concepts from external knowledge graphs for generating more engaging and figurative stories (e.g., the concept of ‘graduation ceremony’ following the detection of an ‘academic gown’ object).
The process of integrating external knowledge has various challenges, ranging from selecting and retrieving relevant knowledge to accurately representing and using it during text generation. Robust retrieval systems that can holistically extract the essence of image sequences or video frames, including the various entities and their interrelationships, are required. Typically, the retrieved knowledge is concatenated with input representations, which are then used for generating text either through fine-tuning Yang et al. (2024) or by prompting general-purpose VLMs Bhattacharyya et al. (2023). However, this approach might lead to models either over- or under-exploiting the retrieved knowledge, potentially leading to incoherent text Gao et al. (2023). To address this, fusion mechanisms that can effectively balance information from representations of both input data and the retrieved knowledge have yet to be developed. Furthermore, retrieving relevant knowledge from increasingly large databases could be computationally expensive, especially in scenarios where visual events unfold over time. Finding methods to optimize retrieval components to improve efficiency is an active research area.
@@
Coherence is the property of text that refers to the ordering of its constituents (words, sentences) and how they relate to each other Althaus, Karamanis, and Koller (2004). Coherent text should have a consistent logical structure in which the events, interactions, and relationships between various elements are ordered in a meaningful way. It is an important aspect of discourse and has been studied extensively in neural language generation Pishdad et al. (2020). For several NLG tasks involving visual events, it might be challenging to ensure that multiple sentences in the generated text are locally coherent. In Movie Auto AD and Visual Storytelling, where there are multiple characters and various interactions unfolding across the sequence of images or video frames, it is often difficult for models to balance between selecting the visual information and representing it cohesively using language Lei et al. (2020). This challenge is more apparent in the Visual Storytelling task, in which models are expected to keep track of multiple things, such as emotional arcs or motivations of the characters and the overarching narrative Surikuchi, Fernández, and Pezzelle (2024). There is increasing work in unimodal text-only storytelling suggesting how using concepts of narratology Piper, So, and Bamman (2021); Antoniak et al. (2024), such as the Genette (1980)’s triangle, can potentially aid models in generating stories with engaging and coherent structures. However, how these theories can be applied to multimodal scenarios where the generated text needs to be consistent with visual input is still overlooked. We argue that this is a promising avenue that should be explored in future work.
@@
Theory of Mind (ToM), which is considered the basis of human social cognition Premack and Woodruff (1978), is described as the ability to understand and make inferences about the mental states (e.g., beliefs, intentions, and desires) of other people or living beings. In the context of the tasks explored in this paper, ToM refers to the ability of models to go beyond merely recognizing actions and to reason about the mental states of entities depicted in the visual input. A recent benchmark for the Video QA task focuses on ToM and shows that existing models lack this skill Mao et al. (2024). We argue that ToM is relevant for all the tasks involving language generation from visual events and that it is closely connected to the other challenges and abilities discussed so far. For instance, in tasks such as Visual Storytelling, to causally connect heterogeneous images in the input sequence, models need to be equipped with different reasoning abilities pertaining to emotions, social perceptions, and intentions. This enables the generation of stories that reflect actions and mental states of the characters beyond the literal interpretation of the visual data.
@@
Recently, several ToM benchmarks have been proposed to assess general-purpose VLMs Gao et al. (2024); Chen et al. (2024b) along different aspects, such as temporal localization of emotions, intentionality-understanding, and perspective-taking. These studies found that only models that are fine-tuned on curated ToM datasets exhibit any reasoning abilities, albeit not aligning with the well-established ToM frameworks explaining human social cognition. Such curated data is scarcely available, and it is currently an open question what alternative architectures or training objectives would enable models to obtain the required ToM abilities. Once again, we claim that future work in this domain should pay more attention to this topic, as mastering ToM will likely lead to models that can better solve tasks involving visual events unfolding over time.
5 Dimensions of Variation
(a) Between consecutive images or video frames of input.
(b) Between consecutive sentences in ground-truth text.
Figure 3: Similarity scores obtained for the tasks along the visual and textual dimensions. We excluded Video QA from the textual analysis due to the lack of multi-sentence datasets.
In the previous section, we discussed several key challenges that apply to the problem space of generating language from visual events. At the same time, different tasks within this space may have different characteristics, due in primis to their various communicative intents. We argue that identifying dimensions of variation across tasks can be valuable to guide the design of training procedures and model architectures, as well as to develop more targeted and appropriate evaluation measures. In this section, we briefly examine two possible dimensions of variation: the similarity of the elements in the visual input, and the consistency of the textual output.
On the visual side, the images or frames in the input can be more or less similar to each other, depending on the task. For instance, in the Change Captioning task, where the goal is to localize and describe changes between a pair of images obtained from real-time surveillance cameras or large-scale remote-sensing snapshots, we expect the similarity between input images to be generally high. On the other hand, tasks as Visual Storytelling, where the images typically depict events loosely forming an overarching narrative, we expect low similarity between consecutive images within each data sample (<visual sequence, text> pair). Regarding the textual output, we similarly expect that the consistency of consecutive sentences within each data sample will be higher or lower depending on the corresponding communicative intent of the task.
We quantitatively analyze the characteristics of five datasets corresponding to the sample tasks considered. Specifically, for each data sample, we compute visual similarity scores between the embeddings of each consecutive image. We use the CLIP visual encoder Radford et al. (2021). Similarly, we compute textual consistency scores between CLIP text encoder embeddings of consecutive sentences in the corresponding ground-truth text of each data sample. For this study, we randomly select 100 instances per task from five datasets—VIST Huang et al. (2016) for Visual Storytelling, Spot-the-diff Jhamtani and Berg-Kirkpatrick (2018) for Change Captioning, Charades Sigurdsson et al. (2016) for Video Captioning, MSVD-QA Xu et al. (2017) for Video QA, and MAD-v1 Soldan et al. (2022) for Movie Auto AD.
Figure 4: Relative positioning of the tasks in the visual-textual shared space, based on the similarity scores obtained for the corresponding datasets. To illustrate, a task in the top right corner of the plot is both highly consistent at the textual level and highly similar at the visual level. Since, for Video-QA, there are no datasets containing multi-sentence datasets, we arbitrarily set its average textual consistency value to $1$ to allow visualization.
This analysis shows that various tasks within the same problem space differ significantly in terms of the two dimensions considered. While this analysis constitutes a small case study, it opens the door to investigating other dimensions that could be relevant to differentiate between tasks even further, and hence to identify key features of concrete tasks. We argue that these and similar analyses should be used to inform modeling and evaluation decisions within the shared space of NLG problems dealing with visual events.
6 Open Problems and Research Directions

### v3

@@ 2.4. Off-the-shelf Pre-trained VLMs
Qwen2.5-VL (Bai et al., 2025) ViT Linear Qwen2.5
DeepSeek-VL (Lu et al., 2024) SigLIP, SAM-B Linear (#2) DeepSeek
LLaVA-Next (Liu et al., 2023) CLIP-ViT-L Linear Mistral
Video-LLaMA (Zhang et al., 2023) ViT-G/14 Q-Former LLAMA
m-PLUG-Owl3 (Ye et al., 2025) SigLIP SO (400M) Linear Qwen2
Video-LLaVA (Lin et al., 2024) OpenCLIP ViT-L Linear (#2) Vicuna v1.5
@@
Table 2. A selection of pre-trained multi-purpose VLMs that are used off-the-shelf for various NLG tasks involving visual events.
3. Properties of Datasets
(a) Between consecutive images or video frames of the visual modality.
(b) Between consecutive sentences in ground-truth text.
[…]
Figure 3. Similarity scores obtained for the tasks along the visual and textual dimensions. We excluded Video QA from the textual analysis due to the lack of multi-sentence datasets.
In the previous section, we summarized modeling approaches and discussed the components of the common architecture used for generating language from visual events. However, different tasks within this space may have different characteristics, due in primis to their various communicative intents. In this section, we briefly examine two possible features of datasets used to train models for the examined tasks: the similarity of the elements in the visual modality and the consistency of the textual modality.
The images or frames on the visual side can be more or less similar to each other, depending on the task. For instance, in the Change Captioning task, where the goal is to localize and describe changes between a pair of images obtained from real-time surveillance cameras or large-scale remote-sensing snapshots, we expect the similarity between images within pairs to be generally high. On the other hand, in tasks such as Visual Storytelling, where the images typically depict events loosely forming an overarching narrative, we expect low similarity between consecutive images within each data sample (¡visual sequence, text¿ pair). Regarding the textual modality, we similarly expect that the consistency of consecutive sentences within each data sample will be higher or lower depending on the corresponding communicative intent of the task. For example, in the Movie Auto AD task, where the language modality is intended to creatively complement the dialogues and soundtrack of the movie, we expect relatively low consistency between consecutive sentences within each data sample.
To illustrate these properties, we quantitatively analyze the characteristics of five datasets corresponding to the sample tasks considered. Specifically, for each data sample, we compute visual similarity by computing cosine similarity scores for each pair of consecutive images. We use the CLIP visual encoder (Radford et al., 2021). Similarly, we operationalize textual consistency by computing cosine similarity scores between CLIP text encoder embeddings of consecutive sentences in the corresponding ground-truth text of each data sample. For this study, we randomly select 100 instances per task from five datasets—VIST (Huang et al., 2016) for Visual Storytelling, Spot-the-diff (Jhamtani and Berg-Kirkpatrick, 2018) for Change Captioning, Charades (Sigurdsson et al., 2016) for Video Captioning, MSVD-QA (Xu et al., 2017) for Video QA, and MAD-v1 (Soldan et al., 2022) for Movie Auto AD.
[…]
Figure 4. Relative positioning of the tasks in the visual-textual shared space, based on the similarity scores obtained for the corresponding datasets. To illustrate, a task in the top right corner of the plot is both highly consistent at the textual level and highly similar at the visual level. Since, for Video-QA, there are no datasets containing multi-sentence datasets, we arbitrarily set its average textual consistency value to 11 to allow visualization.
This analysis shows that properties of data may differ significantly across tasks within the same problem space. While this analysis constitutes a small case study, we hope that it will encourage the community to explore the tasks further for identifying other features of the data that researchers aim to model. We believe that these and similar analyses of datasets have several benefits. First, they can be used to inform modeling and evaluation decisions within the shared space of NLG problems dealing with visual events. In the next section, we specifically discuss current approaches to evaluation. Furthermore, understanding the communicative intents behind specific tasks and analyzing properties of their data can modulate the importance of different abilities required for approaching them optimally. In Section 5, we identify key open challenges related to such abilities that are common across tasks and describe them in a unified manner.
4. Evaluation Approaches
@@
Automatically generated text is notoriously difficult to evaluate. Most of the methods currently used to evaluate text describing visual events are borrowed from simpler NLG scenarios, and hence are not sufficient to fully understand the quality of language generated from visual input unfolding over time. These methods range from using traditional n-gram matching metrics to obtaining human judgments and ratings to, more recently, using off-the-shelf pre-trained VLMs assessing the generated output. In this section, we broadly classify evaluation methods into two main categories—automatic and human evaluation—and discuss several metrics widely used for vision-to-text tasks, along with the rationales for using them. Through this discussion, we also underline how the current approaches are lacking when it comes to assessing models with respect to the data features and the particular challenges relevant to the problem of language generation from visual events.
4.1. Automatic Evaluation
@@
Very often, tasks involving language generation from visual events assess a model-generated candidate text by comparing it to corresponding human-written references. Specifically, traditional metrics that were originally designed for evaluating machine translation and text summarization tasks—BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and ROUGE (Lin, 2004)—are used to measure precision and recall of overlapping $\mathit{n}$-grams between the candidate and a reference text. Usually, metrics such as CIDEr (Vedantam et al., 2015) and SPICE (Anderson et al., 2016), which have been specifically developed for the evaluation of image and video captioning, are also used in conjunction with the above three metrics.
All the above-mentioned metrics rely on direct raw text comparisons of ground-truth references and model outputs. As this ground-truth data might not be always available, embedding-level reference-based evaluation metrics such as WMD (Kusner et al., 2015), BERTScore (Zhang* et al., 2020), and ViLBERTScore (Lee et al., 2020) have been proposed. Recent work for Video Captioning, Video QA, and Movie Auto AD have used these metrics to measure the similarity between candidate and reference embeddings, obtained by projecting corresponding text into a common pre-trained semantic space (Lin et al., 2022; Han et al., 2023a, 2024).
@@
Comparing model-generated text to ground-truth references, typically provided by crowd-workers or scraped from the internet, has various limitations when applied to the context of visually-grounded language generation. First, most reference-based metrics do not account for the visual modality upon which the generated text is conditioned. Second, reference-based metrics are generally inappropriate to accurately evaluate language that is not merely descriptive, but may involve creative or narrative concepts. This is the case for various tasks involving visual events, such as Video Captioning, Movie Auto AD, and Visual Storytelling, where the data shows a low degree of textual consistency between consecutive sentences, as shown in Figure 4.
On the language-only level, various reference-free metrics—not relying on any reference text—such as MAUVE (Pillutla et al., 2021) and UNION (Guan and Huang, 2020) have been proposed. However, for visually-conditioned text generation tasks, this shift is still relatively recent, with a persistent emphasis on reference-based $\mathit{n}$-gram metrics to date. That said, various reference-free metrics have been recently proposed to assess different aspects of evaluation that are important for generation tasks that specifically concern visual events. For instance, metrics such as CLIPScore (Hessel et al., 2021) and GROOViST (Surikuchi et al., 2023) have been developed for evaluating visual grounding—the degree of alignment between the generated text and the visual input—in Video Captioning and Visual Storytelling tasks. Similarly, the RoViST (Wang et al., 2022) suite of metrics has been proposed to assess coherence, the extent of repetition, and visual grounding of the generated text. Yet other metrics, such as CRITIC (Han et al., 2024) and ‘character matching’ (Hong et al., 2023), are designed to evaluate task-specific aspects such as the accuracy of referencing to characters in the model outputs.
Finally, there is currently an increasing adoption and reliance on using pre-trained LLMs and VLMs as judges (Li et al., 2024b). Essentially, these pre-trained general-purpose models are prompted to score or rate a model-generated response along any of the evaluation dimensions of interest, such as fluency or relevance (e.g., ‘How fluent is the generated text on a scale of 1 to 5?’). However, the effectiveness and reliability of this approach are still debated (Bavaresco et al., 2024).
4.2. Human Evaluation
@@
Participants of these evaluation efforts are provided with a set of task-specific rubrics along with representative examples required for judging the model outputs. They are asked to assess the overall quality of model-generated outputs either independently (per sample) (Surikuchi et al., 2023) or relative to outputs from other models (Wang et al., 2022). Alternatively, evaluators might be required to provide ratings for various criteria ranging from broad (e.g., text conciseness, fluency, grammatical correctness) to specific (e.g., factuality, hallucinations, expressiveness). The obtained scores are usually compared pairwise to rank models appropriately.
Some pre-trained VLM frameworks, such as LLaVA-RLHF (Sun et al., 2024) leverage this qualitative feedback to optimize model parameters for learning to generate human-preferred text. Although human evaluation is still indispensable for several tasks, it is also expensive, time-consuming, and challenging. Defining clear evaluation protocols for ensuring the reliability and quality of human judgments is an active research area (Kasai et al., 2022; Ruan et al., 2024).
@@
5. Key Challenges in Language Generation from Visual Events
As mentioned in the Introduction, tasks involving language generation from visual events present a unique set of challenges. These involve modeling the intricate relationship between temporally ordered visual events and the structure, content, and function of the language used to interpret, describe, or narrate them. We argue that solving these tasks requires models to be capable of identifying and managing such intricacies, leveraging the temporal, causal, and discourse-level patterns inherent in both visual data and language, as well as the many ways in which these aspects interact. In this section, we identify and describe five main open challenges and discuss how they are key to various tasks. This unified perspective is a novel contribution of our work, which we believe can help the community to find common strategies for making progress across tasks.
@@ 5. Key Challenges in Language Generation from Visual Events
Identifying and tracking entities is an important requisite for accurately interpreting actions and relationships between them. This has been underlined by several works in both language understanding (Paperno et al., 2016; Kim and Schuster, 2023) and computer vision domains (Narayan et al., 2017; Luo et al., 2021). In our multimodal tasks, the availability of input signals from two modalities makes the aspect of disambiguating entities more challenging. For instance, in the Movie Auto AD or Visual Storytelling tasks where the visual input is heterogeneous, entities in the visual input tend to ‘disappear’ in some of the images or video frames at the intermediate temporal positions, while still being actively referenced in the textual input at the corresponding positions. Surikuchi et al. (2023) have denoted such cases as being temporally misaligned. To track entities accurately under such temporal misalignment, it is necessary to not only learn causal relations between people and objects in each of the modalities at corresponding positions, but also to obtain a cross-modal, cross-temporal representation of all the relationships relevant to the scene and the overarching narrative. For the Change Captioning, Video Captioning, or Video QA tasks, in which the input images or video frames are typically similar to each other, it is crucial to differentiate meaningful semantic entities and their changes from various distractions. While in Change Captioning, viewpoint changes or illuminations are considered as distractors and discarded, in the Video QA task, entities relevant to answering the question need to be differentiated from others for accurate tracking. Effective tracking of entities would therefore require accounting for changes in appearance (including disappearance), capturing interactions, and correctly identifying occlusions.
@@
Humans acquire language through perception and interaction with the environment (Barsalou, 2008; Iverson and Goldin-Meadow, 2005) and consequently, this enables them to seamlessly ground language in visual data. Over the years, a great deal of work has been proposed to adapt the architecture and learning process of vision-language models for acquiring visual grounding (Suglia et al., 2024). However, there are still significant challenges for achieving human-levels of grounding using computational models, and this becomes more apparent when looked at from the perspective of the tasks under investigation here. For instance, in Visual Storytelling or Movie Auto AD, the language used is often creative or abstract, including various sources of underspecification, such as context-dependent adverbs like ‘there’, ‘often’, or ‘today’ (’a photographer over there clicks…’). Recently, vision-language models have been shown to struggle with underspecified language, due to the challenges related to grounding underspecified terms to regions in the visual content (Pezzelle, 2023). Moreover, the amount of language informativeness—i.e., the degree of information required for identifying the correct object (Coppock et al., 2020)—could be inadequate in tasks such as Video QA, particularly with the presence of confounding entities in various frames (e.g., input video of a football match and a question: ‘What is the color of the card the referee is holding?’). In addition, when grounding objects, models have often struggled to reliably capture spatial relationships (Kamath et al., 2023). To summarize, the language used in these tasks might not be merely descriptive but rather complementary to the content depicted in the corresponding images or videos. This makes visual grounding challenging without access to relevant additional external knowledge.
@@
To meet the communicative specificity of each task dealing with visual events, models would need to use additional information beyond what is available in the visual input. When performing Video Captioning or Video QA on specific domains such as news, the input video might not contain all the aspects needed to correctly describe its contents or answer questions about it (Whitehead et al., 2018; Jin et al., 2023). To address this, various approaches often rely on using large pre-trained general purpose models or external knowledge bases such as ConceptNet (Speer et al., 2017) to retrieve both factual and commonsense information. This method is commonly referred to as retrieval-augmented generation (RAG) (Lewis et al., 2020). Besides being sources for missing information, external knowledge bases are also leveraged for enriching the generated text with social or cultural contexts. For instance, in Visual Storytelling, some approaches use recognized visual objects in the input to retrieve concepts from external knowledge graphs for generating more engaging and figurative stories (e.g., the concept of ‘graduation ceremony’ following the detection of an ‘academic gown’ object).
The process of integrating external knowledge has various challenges, ranging from selecting and retrieving relevant knowledge to accurately representing and using it during text generation. Robust retrieval systems that can holistically extract the essence of image sequences or video frames, including the various entities and their interrelationships, are required. Typically, the retrieved knowledge is concatenated with input representations, which are then used for generating text either through fine-tuning (Yang et al., 2024) or by prompting general-purpose VLMs (Bhattacharyya et al., 2023). However, this approach might lead to models either over- or under-exploiting the retrieved knowledge, potentially leading to incoherent text (Gao et al., 2023). To address this, fusion mechanisms that can effectively balance information from representations of both input data and the retrieved knowledge have yet to be developed. Furthermore, retrieving relevant knowledge from increasingly large databases could be computationally expensive, especially in scenarios where visual events unfold over time. Finding methods to optimize retrieval components to improve efficiency is an active research area.
@@
Coherence is the property of text that refers to the ordering of its constituents (words, sentences) and how they relate to each other (Althaus et al., 2004). Coherent text should have a consistent logical structure in which the events, interactions, and relationships between various elements are ordered in a meaningful way. It is an important aspect of discourse and has been studied extensively in neural language generation (Pishdad et al., 2020). For several NLG tasks involving visual events, it might be challenging to ensure that multiple sentences in the generated text are locally coherent. In Movie Auto AD and Visual Storytelling, where there are multiple characters and various interactions unfolding across the sequence of images or video frames, it is often difficult for models to balance between selecting the visual information and representing it cohesively using language (Lei et al., 2020). This challenge is more apparent in the Visual Storytelling task, in which models are expected to keep track of multiple things, such as emotional arcs or motivations of the characters and the overarching narrative (Surikuchi et al., 2024). In such tasks, there is a tension between coherence and creativity (cf. low textual consistency of Visual Storytelling data in Figure 4). There is increasing work in unimodal text-only storytelling suggesting how using concepts of narratology (Piper et al., 2021; Antoniak et al., 2024), such as the Genette (1980)’s triangle, can potentially aid models in generating stories with engaging and coherent structures. However, how these theories can be applied to multimodal scenarios where the generated text needs to be consistent with visual input is still overlooked. We argue that this is a promising avenue that should be explored in future work.
@@
Theory of Mind (ToM), which is considered the basis of human social cognition (Premack and Woodruff, 1978), is described as the ability to understand and make inferences about the mental states (e.g., beliefs, intentions, and desires) of other people or living beings. In the context of the tasks explored in this paper, ToM refers to the ability of models to go beyond merely recognizing actions and to reason about the mental states of entities depicted in the visual input. A recent benchmark for the Video QA task focuses on ToM and shows that existing models lack this skill (Mao et al., 2024). We argue that ToM is relevant for all the tasks involving language generation from visual events and that it is closely connected to the other challenges and abilities discussed so far. For instance, in tasks such as Visual Storytelling, to causally connect heterogeneous images in the input sequence, models need to be equipped with different reasoning abilities pertaining to emotions, social perceptions, and intentions. This enables the generation of stories that reflect actions and mental states of the characters beyond the literal interpretation of the visual data.
@@
Recently, several ToM benchmarks have been proposed to assess general-purpose VLMs (Gao et al., 2024; Chen et al., 2024b) along different aspects, such as temporal localization of emotions, intentionality-understanding, and perspective-taking. These studies found that only models that are fine-tuned on curated ToM datasets exhibit any reasoning abilities, albeit not aligning with the well-established ToM frameworks explaining human social cognition. Such curated data is scarcely available, and it is currently an open question what alternative architectures or training objectives would enable models to obtain the required ToM abilities. Once again, we claim that future work in this domain should pay more attention to this topic, as mastering ToM will likely lead to models that can better solve tasks involving visual events unfolding over time.
6. Open Problems and Research Directions